OpenStudy (anonymous):

(5-x)^1/2 = x+1 (show steps plz!)

OpenStudy (anonymous):

How about I take you through it step by step instead?

OpenStudy (anonymous):

sounds good to me.

OpenStudy (anonymous):

Okay, so it's easier if you get rid of the 1/2 exponent, so square both sides.

OpenStudy (anonymous):

i did that and got 5-x= x^2+2x+1

OpenStudy (anonymous):

then set everything equal to 0

OpenStudy (anonymous):

0=x^2+3x-4

OpenStudy (anonymous):

factor that

OpenStudy (anonymous):

everything after is a blurr

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ummm yeah. i have no idea.

OpenStudy (anonymous):

Quadratic formula....

OpenStudy (anonymous):

all right. so when you factor you're going to get (x+y)(x-z)=0 where y and z are numbers, possibly the same number. You know you're going to subtract one of them and add one of them because if you look at the number that has no x's in it (-4) you subtract it. Did that much make sense?

OpenStudy (anonymous):

You can also use the quadratic formula, but I find factoring it easier.

OpenStudy (anonymous):

Factoring is easier, but for people who can't do that sort of factoring in their head, the quadratic formula is handy.

OpenStudy (anonymous):

Or at least it's easier if you understand factoring, which is definitely something you learn by practicing.

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

is that -(b)sqrt(b^2)-4(a)(c)/ 2(a)?

OpenStudy (anonymous):

That's true, I completely agree with you, Harwin. Sorry, slow typer apparently!

OpenStudy (anonymous):

(-b + sqrt(b^2 - 4ac))/2a and (-b - sqrt(b^2 - 4ac))/2a

OpenStudy (anonymous):

oh whoops yeah for got the +-

OpenStudy (anonymous):

Anyway, snooozzzzze knows how to get you through it. I'll give the show back to him.

OpenStudy (anonymous):

haha her, but that's ok. :) Do you want to use the quadractic formula, or try factoring?

OpenStudy (anonymous):

so does \[-11+-\sqrt{73}/8\] sound like its on the right track?

OpenStudy (anonymous):

Oh sorry. English really needs a genderless pronoun other than "it"....

OpenStudy (anonymous):

ha

OpenStudy (anonymous):

Use the quadractic formula on this equation: 0=x^2+3x-4

OpenStudy (anonymous):

oh shoot sorry i was looking at another equation

OpenStudy (anonymous):

That's very true. Not sure calling me it would have been better, but in any case it's not a big deal at all. RedCarr: No problem at all. Just try again with the right equation this time.

OpenStudy (anonymous):

how about \[-3+-\sqrt{57}/2\]?

OpenStudy (anonymous):

I'm not sure how you got the 57, but it's wrong.

OpenStudy (anonymous):

also, when you do it, make sure that you know that the 2 is over everything, not just the squareroot part.

OpenStudy (anonymous):

in place of the 57 would it be 25?

OpenStudy (anonymous):

that looks right to me.

OpenStudy (anonymous):

so x=1 and X=-8?

OpenStudy (anonymous):

No, this is what you should get: (−3+5)/2 (−3-5)/2

OpenStudy (anonymous):

? i meant -4 but why the parenthesis?

OpenStudy (anonymous):

because otherwise you'd have written -3+5/2 which would simplify to -3+2.5=-0.5

OpenStudy (anonymous):

but yes, -4 and 1 is correct.

OpenStudy (anonymous):

except look at your original problem to see if they work.

OpenStudy (anonymous):

do you have a graphing calc with you? if you do plug in \[-3+\sqrt{25}\]

OpenStudy (anonymous):

because actually only one of them works in that case.

OpenStudy (anonymous):

I don't have a calculator, but I can visualize it.

OpenStudy (anonymous):

yeah only 1 does not the -4

OpenStudy (anonymous):

right, and that's because when we started the problem we squared both sides, which gave us the extra answer.

OpenStudy (anonymous):

well when i plugged it in my answer was 2. then i divided that 2 by 2 to get 1 which is the only answer.

OpenStudy (anonymous):

I'm sorry, could you restate that? I'm not really sure what you mean.